Determine whether each pair is parallel, perpendicular, or neither.
x-4y= -12 y=4x-4

Answer by shree840(260)

Put the equations in the form of y=mx+b
Parallel line will have same m
perpendicular will have m1*m2=-1
1. x-4y=-12
4y=x+12
y=x/4+3 for perpendicular flip the m and change the sign 1/4 becomes -4
y=-4x+b
so lines are neithe parallel or perpendicular.
draw the lines on ti84